# Weighted shift operators on spaces of analytic functions over an annulus - I

### v1

@@ 4. Adjoints of unilateral weighted shifts
In this section, we study the dynamical properties of the adjoint of the unilateral forward shift $F_{w}(\sum_{n=0}^{\infty}\lambda_{n}z^{n})=\sum_{n=0}^{\infty}\lambda_{n}w_{n}z^{n+1}$ defined on $\ell^{p}_{a,b}(\mathbb{D})$, where $w=\{w_{n}\}_{n=0}^{\infty}$ is a sequence of non-zero complex numbers. The differences between $F_{w}$ acting on $\ell^{p}_{a,b}(\mathbb{D})$ and the classical weighted forward shifts on $\ell^{p}$ will be pointed out in Section $5$.
First, we derive some sufficient conditions for the unilateral $F_{w}$ to bounded in terms of $\{a_{n}\}$, $\{b_{n}\}$ and the weight sequence $\{w_{n}\}$. For the unweighted right right $F$, some necessary and sufficient conditions for $F$ to be bounded on $\ell^{2}_{a,b}(\mathbb{D})$ were obtained by Adams and McGuire [1]. Weighted backward shifts $B_{w}$ on $\ell^{p}_{a,b}(\mathbb{D})$ were studied by Das and Mundayadan [15].
@@
where the coefficient for $j=0$ should be understood as $1$. (We refer to [1] for the previous calculations in the context of the study of tridiagonal kernels. Also, see [15].)
Acting $F_{w}$ on the normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\geq 0$ in $\ell^{p}_{a,b}(\mathbb{D})$, we obtain
$F_{w}(f_{n})(z)$ $=$ $w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}$
$=$ $\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+\left(\frac{w_{n+1}b_{n}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right)a_{n+2}z^{n+2}$
$=$ $\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+c_{n}\sum_{j=0}^{\infty}(-1)^{j}\left(\frac{\prod_{k=0}^{j-1}b_{n+2+k}}{\prod_{k=0}^{j-1}a_{n+3+k}}\right)f_{n+2+j},$
where
$c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}.$
@@
In the following theorem we provide a sufficient condition for $F_{w}$ to be bounded. For the particular cases of $p=2$ and the unweighted forward shift, we refer to [1], [16], and [17].
@@
$\sup_{n\geq 0}~{}\Big{\{}\left\lvert\frac{w_{n}a_{n}}{a_{n+1}}\right\rvert,~{}|c_{n}|\Big{\}}<\infty,~{}\text{and}~{}\sum_{n=3}^{\infty}\sup_{j\geq 0}~{}\left|\frac{c_{j}b_{j+2}b_{j+3}\cdots b_{j+n-1}}{a_{j+3}a_{j+4}\cdots a_{j+n}}\right|<\infty,$
then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$, where $1\leq p<\infty$.
@@
We give only an outline of the proof. The matrix of $F_{w}$ can be formally written as $[F_{\alpha}]$+$\sum_{n=2}^{\infty}[F_{n}]$. Here, $[F_{\alpha}]$ is the matrix of the standard weighted forward shift $F_{\alpha}(e_{i})\mapsto\alpha_{i}e_{i+1}$ on $\ell^{p}$, $i\geq 1$, having weights
@@
on $\ell^{p}$. The matrix $[F_{n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th subdiagonal, where $n\geq 2$. Observe that $[F_{n}]$ is the matrix of suitable powers of a weighted forward shift $F_{n}$ for $n\geq 2$.
It follows, respectively by the first two assumptions in the theorem, that the weighted shift $F_{\alpha}$ and $F_{2}$ are bounded on $\ell^{p}$. Since
@@
the third condition in the theorem gives that $F_{n}$ is bounded, and $\sum_{n\geq 3}\|F_{n}\|<\infty$, with respect to the operator norm. Hence, the shift $F_{w}$ is bounded. This completes the proof of the theorem. ∎
@@
Let $X$ be a Banach space of analytic functions on the unit disc, having evaluation functionals continuous. Let $k_{n}$, $n\geq 0$, denote the $n$-th Taylor coefficient functional. Then, the span of $\{k_{n}:n\geq 0\}$ is weak-dense in $X^{*}$, and norm dense if $X^{*}$ is reflexive.
Using above result, we prove:
@@
Let $\ell^{p}_{a,b}(\mathbb{D}),$ $1<p<\infty,$ be the analytic Banach space over the unit disc $\mathbb{D}$ corresponding to $a=\{a_{n}\}_{n=0}^{\infty}$ and $b=\{b_{n}\}_{n=0}^{\infty}$ satisfying the conditions in (4.2). Then, the following hold for the adjoint $F_{w}^{*}$:
@@
    $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathbb{D})^{*}$ if and only if
    $\liminf_{n\rightarrow\infty}\Big{|}\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\Big{|}=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (4.3)
@@
    $F_{w}^{*}$ is topologically mixing if and only if
    $\lim_{n\rightarrow\infty}\Big{\lvert}\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\Big{\rvert}=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (4.4)
@@
(i) First suppose that (4.3) holds, and we will use Gethner-shapiro criterion to show that $F_{w}^{*}$ is hypercyclic. Consider $D:=\text{span}\{k_{n}:n\geq 0\}$. Then $D$ is dense in $\ell^{p}_{a,b}(\mathbb{D})^{*},$ by the previous proposition. Define $S:D\rightarrow D$ by
$S(k_{n})=\frac{1}{w_{n}}k_{n+1},$
$\forall~{}~{}n\geq 0.$ Since $F_{w}^{*}S(f)=f$ and $(F_{w}^{*})^{n}f\to 0,$ as $n\to\infty,$ $\forall$ $f\in D,$ we just need to show that $S^{n_{k}}(k_{\nu})$ tends to 00, as $k\to\infty,$ for any increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ But this is clear since by Proposition (2.2) and $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
for some constants $M_{1}>0.$ Now by the assumption (4.3), we get that $F_{w}^{*}$ satisfies the Gethner-Shapiro criterion. Hence, $F_{w}^{*}$ is hypercyclic.
Conversely, assume that $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathbb{D})^{*}$ and we have to show that (4.3) holds. Let $f=z^{\nu},\nu\in\mathbb{N}$ and $g$ be a non-zero vector in $\ell^{p}_{a,b}(\mathbb{D})^{*}.$ Then
$\langle(F_{w}^{*})^{n}g,f\rangle=\langle g,F_{w}^{n}f\rangle=\langle g,w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}\rangle.$
Now by Proposition 4.1 of [15] we know that
$\lVert w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}\rVert\leq M\left\lvert\frac{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}{a_{\nu+n}}\right\rvert,$
for some constant $M>0.$ If possible, let $\lim\inf_{n\rightarrow\infty}\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert>0,$ then we get that $\{(F_{w}^{*})^{n}g\}_{n\geq 0}$ is a bounded set. This contradicts the fact that $F_{w}^{*}$ has a hyperyclic vector, and hence, $\lim\inf_{n\rightarrow\infty}\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert=0.$
(ii) Apply the Gethner-Shapiro criterion with $n_{k}=k$ for $k\geq 1$, and proceed as above.
(iii) The supercyclicity criterion with $D$ and $S$ as above, taking $n_{k}=k$, $k\geq 1$, is satisfied.
Our next result on the chaos of $F_{w}^{*}$. In [33], the authors provided a sufficient condition for the adjoint $F^{*}$ to be chaotic on a general analytic reproducing kernel space over $\mathbb{D}$ which includes the case $\ell^{2}_{a,b}(\mathbb{D})$. The following representations of $f_{n}^{*}$ in terms of $k_{n}$ will be utilized to get the chaos of the adjoint of forward weighted shifts.
@@
Let $k_{n}$ denote the co-ordinate functional
$f\mapsto\frac{f^{(n)}(0)}{n!},$
defined on $\ell^{p}_{a,b}(\mathbb{D})$ or $c_{0,a,b}(\mathbb{D})$, where $n\geq 0$. We have
$f_{n}^{*}=\frac{1}{a_{n}}k_{n}-\frac{b_{n-1}}{a_{n}a_{n-1}}k_{n-1}+\frac{b_{n-1}b_{n-2}}{a_{n}a_{n-1}a_{n-2}}k_{n-2}-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}k_{0},$
@@
Proof.
If $f\in\ell^{p}_{a,b},$ then $f(z)=\sum_{n=0}^{\infty}\lambda_{n}f_{n}(z)=\sum_{n=0}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}.$ Rearranging into a power series, we obtain $f(z)=\lambda_{0}a_{0}+\sum_{n\geq 1}(\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1})z^{n},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}$ and $k_{0}(f)=\lambda_{0}a_{0}.$ Now for $f\in\ell^{p}_{a,b},$ we have
$\frac{1}{a_{n}}k_{n}(f)-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}k_{0}(f)$ $=$ $\frac{1}{a_{n}}(\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1})-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}(\lambda_{0}a_{0})$
$=$ $\lambda_{n}$
$=$ $f_{n}^{*}(\sum_{m=0}^{\infty}\lambda_{m}f_{m})$
$=$ $f_{n}^{*}(f),$
for all $n\geq 1$ and $\frac{1}{a_{0}}k_{0}(f)=\frac{1}{a_{0}}(\lambda_{0}a_{0})=\lambda_{0}=f_{0}^{*}(\sum_{m=0}^{\infty}\lambda_{m}f_{m})=f_{0}^{*}(f).$
We now have necessary and sufficient conditions for $F_{w}^{*}$ to be chaotic.
@@
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty.$
@@
$\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{b_{n}b_{n+1}\cdots b_{n+j-1}}{a_{n+1}a_{n+2}\cdots a_{n+j}}\right\rvert^{q}\right)^{\frac{p}{q}}<\infty.$
@@
To see that (ii) implies (iii), let $F=\sum_{m=0}^{\infty}\lambda_{m}f_{m}^{*}$ be a non-zero periodic vector for $F_{w}^{*}$ on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$, where $f_{m}^{*}$, $m\geq 0$, forms an unconditional Schauder basis for $\ell^{p}_{a,b}(\mathbb{D})^{*}$. Now by Proposition 4.5, each $f_{m}^{*},m\geq 1$ can be written as
$f_{m}^{*}=\frac{1}{a_{m}}k_{m}-\frac{b_{m-1}}{a_{m}a_{m-1}}k_{m-1}+\frac{b_{m-1}b_{m-2}}{a_{m}a_{m-1}a_{m-2}}k_{m-2}-\cdots+(-1)^{m}\frac{b_{m-1}\cdots b_{0}}{a_{m}\cdots a_{0}}k_{0},$
@@
and $f_{0}^{*}=\frac{1}{a_{0}}k_{0}.$ Let $\nu\in\mathbb{N}$ such that $(F_{w}^{*})^{\nu}F=F$ for $z\in\mathbb{D}$. Then $(F_{w}^{*})^{n\nu}F=F$ for all $n\geq 1$. Equating the coefficient of $k_{0}$ both sides, we get
$\frac{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}{a_{n\nu}}\left(\lambda_{n\nu}-\lambda_{n\nu+1}\frac{b_{n\nu}}{a_{n\nu+1}}+\cdots\right)=\frac{1}{a_{0}}\left(\lambda_{0}-\lambda_{1}\frac{b_{0}}{a_{1}}+\lambda_{2}\frac{b_{0}b_{1}}{a_{1}a_{2}}-\cdots\right),$
implies that
@@
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{p}\left\lvert\lambda_{0}-\lambda_{1}\frac{b_{0}}{a_{1}}+\lambda_{2}\frac{b_{0}b_{1}}{a_{1}a_{2}}-\cdots\right\rvert^{p}=\left\lvert a_{0}\right\rvert^{p}\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}-\lambda_{n\nu+1}\frac{b_{n\nu}}{a_{n\nu+1}}+\cdots\right\rvert^{p}.$
Using Hölder inequality, we can get
@@
$\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}-\lambda_{n\nu+1}\frac{b_{n\nu}}{a_{n\nu+1}}+\lambda_{n\nu+2}\frac{b_{n\nu}b_{n\nu+1}}{a_{n\nu+1}a_{n\nu+2}}-\cdots\right\rvert^{p}<\infty,$
@@
as $\{\lambda_{n}\}_{n}\in\ell^{p}(\mathbb{N})$ and by assumption $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{b_{n}b_{n+1}\cdots b_{n+j-1}}{a_{n+1}a_{n+2}\cdots a_{n+j}}\right\rvert^{q}\right)^{\frac{p}{q}}<\infty.$ Therefore
@@
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{p}<\infty.$ Similarly, equating the coefficient of $k_{j},j=1,2,\cdots,\nu-1$ both sides we get $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu+j}}{w_{n\nu+j-1}w_{n\nu+j-2}\cdots w_{j}}\right\rvert^{p}<\infty.$ Consequently, the series $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty.$

### v2

@@ 4. Adjoints of unilateral weighted shifts
In this section, we study the dynamical properties of the adjoint of the unilateral forward shift $F_{w}(\sum_{n=0}^{\infty}\lambda_{n}z^{n})=\sum_{n=0}^{\infty}\lambda_{n}w_{n}z^{n+1}$ defined on $\ell^{p}_{a,b}(\mathbb{D})$, where $w=\{w_{n}\}_{n=0}^{\infty}$ is a sequence of non-zero complex numbers. The differences between $F_{w}$ acting on $\ell^{p}_{a,b}(\mathbb{D})$ and the classical weighted forward shifts on $\ell^{p}$ will be pointed out in Section $5$. First, we derive some sufficient conditions for the unilateral $F_{w}$ to bounded in terms of $\{a_{n}\}$, $\{b_{n}\}$ and the weight sequence $\{w_{n}\}$. For the unweighted forward shift $F$, some necessary and sufficient conditions for $F$ to be bounded on $\ell^{2}_{a,b}(\mathbb{D})$ were obtained by Adams and McGuire [2]. Weighted backward shifts $B_{w}$ on $\ell^{p}_{a,b}(\mathbb{D})$ were studied by the authors in [13]. For operator theoretic results in the $p=2$ case, see [14], and [15].
@@
where the coefficient for $j=0$ should be understood as $1$. (See, also, [2].) Note that $z^{n}\in\ell^{p}_{a,b}(\mathbb{D})$ if and only if
$\|z^{n}\|^{p}:=\frac{1}{|a_{n}|^{p}}\sum_{j=0}^{\infty}\left|\frac{\prod_{k=0}^{j-1}b_{n+k}}{\prod_{k=0}^{j-1}a_{n+k+1}}\right|^{p}<\infty.$ (4.2)
Acting $F_{w}$ on the normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\geq 0$, of $\ell^{p}_{a,b}(\mathbb{D})$, we obtain
$F_{w}(f_{n})(z)$ $=$ $w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+\left(\frac{w_{n+1}b_{n}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right)a_{n+2}z^{n+2}$
$\hskip 48.36958pt=\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+c_{n}\sum_{j=0}^{\infty}(-1)^{j}\left(\frac{\prod_{k=0}^{j-1}b_{n+2+k}}{\prod_{k=0}^{j-1}a_{n+3+k}}\right)f_{n+2+j},$
where $c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}$.
@@
We now provide a sufficient condition for $F_{w}$ to be bounded. Only an outline of the proof will be given. For the particular cases of $p=2$ and the unweighted forward shift $F$, we refer to [2].
@@
$\beta_{1}:=\sup_{n\geq 0}~{}\Big{\{}\left\lvert\frac{w_{n}a_{n}}{a_{n+1}}\right\rvert,~{}|c_{n}|\Big{\}}<\infty,~{}\text{and}~{}\beta_{2}:=\sum_{n=3}^{\infty}\sup_{j\geq 0}~{}\left|c_{j}\frac{\prod_{k=3}^{n}b_{j+k-1}}{\prod_{k=3}^{n}a_{j+k}}\right|<\infty,$
then $F_{w}$ is a bounded operator on $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$, where $1\leq p<\infty$. Moreover, $\|F_{w}\|\leq\beta_{1}+\beta_{2}$.
@@
The matrix of $F_{w}$ can be formally written as $[F_{\alpha}]$+$\sum_{n=2}^{\infty}[F_{n}]$. Here, $[F_{\alpha}]$ is the matrix of the standard weighted forward shift $F_{\alpha}(e_{i})\mapsto\alpha_{i}e_{i+1}$ on $\ell^{p}$, $i\geq 1$, having weights
@@
on $\ell^{p}$. Note that the matrix $[F_{n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th subdiagonal, where $n\geq 2$. Observe that $[F_{n}]$ is the matrix of suitable powers of a weighted forward shift $F_{n}$ for $n\geq 2$.
It follows, by the first assumption in the theorem, that the weighted shifts $F_{\alpha}$ and $F_{2}$ are bounded on $\ell^{p}$. Since
@@
the second condition in the theorem gives that $F_{n}$ is bounded, and $\sum_{n\geq 3}\|F_{n}\|<\infty$, with respect to the operator norm. Hence, the shift $F_{w}$ is bounded. This completes the proof of the theorem. ∎
@@
Let $\mathcal{E}$ be a Banach space of analytic functions on the unit disc, having evaluation functionals continuous. Let $k_{n}$, $n\geq 0$, denote the $n$-th Taylor coefficient functional. Then the span of $\{k_{n}:n\geq 0\}$ is weak-dense in $\mathcal{E}^{*}$, and norm dense if $\mathcal{E}$ is reflexive.
In the theorem below, we state the results for hypercyclicity only. For mixing of $F_{w}^{*}$, “liminf” is replaced by “lim”, and we do not state it.
@@
Let $\ell^{p}_{a,b}(\mathbb{D}),$ $1<p<\infty,$ be the analytic Banach space over the unit disc $\mathbb{D}$ corresponding to $a=\{a_{n}\}_{n=0}^{\infty}$ and $b=\{b_{n}\}_{n=0}^{\infty}$. Then, the following hold for the adjoint $F_{w}^{*}$, assuming that $F_{w}$ is bounded.
@@
    If
    $\liminf_{n\rightarrow\infty}~{}\frac{|a_{\nu+n}|+|b_{\nu+n-1}|}{\prod_{k=0}^{n-1}|w_{\nu+k}|}=0,~{}\forall~{}\nu\geq 0,$ (4.4)
    then $F_{w}^{*}$ is hypercyclic.
@@
    Conversely, if $F_{w}^{*}$ is hypercyclic, then
    $\sup_{n\geq 0}~{}\frac{\prod_{k=0}^{n-1}|w_{k}|}{|a_{n}|}\left(\sum_{j=0}^{\infty}\left|\frac{\prod_{k=0}^{j-1}b_{n+k}}{\prod_{k=0}^{j-1}a_{n+k+1}}\right|^{p}\right)^{1/p}=\infty.$
@@
  4. (iv)
    If the conditions in (4.3) are satisfied, then $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$ if and only if
    $\liminf_{n\rightarrow\infty}\left|\frac{a_{\nu+n}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right|=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (4.5)
@@
(i) First suppose that (4.4) holds, and we will use Gethner-shapiro criterion to show that $F_{w}^{*}$ is hypercyclic. Consider $D:=\text{span}\{k_{\nu}:\nu\geq 0\}$. Then $D$ is dense in $(\ell^{p}_{a,b}(\mathbb{D}))^{*},$ by the previous proposition. Define $S:D\rightarrow D$ by
$S(k_{\nu})=\frac{1}{w_{\nu}}k_{\nu+1},$
$\forall~{}~{}\nu\geq 0.$ Since $F_{w}^{*}S(f)=f$ and $(F_{w}^{*})^{n}f\to 0,$ as $n\to\infty,$ $\forall$ $f\in D,$ we just need to show that there exists $(n_{k})$ such that $S^{n_{k}}(k_{\nu})$ tends to 00, as $k\to\infty,$ for $\nu\geq 0$. Indeed, Proposition 2.2 along with our assumption in (i) gives that, for some $(n_{k})$,
$\lVert S^{n_{k}}(k_{\nu})\rVert=\left\lVert\frac{k_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rVert\leq M_{1}\frac{|a_{\nu+n_{k}|}+|b_{\nu+n_{k}-1}|}{|w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}|}\longrightarrow 0,$
as $k\rightarrow\infty$, where $M_{1}$ is a constant. Thus, $F_{w}^{*}$ satisfies the Gethner-Shapiro criterion.
Conversely, if $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$, then by the previously mentioned result of Bonet, we have
$\sup_{n\geq 1}\|w_{0}w_{1}\cdots w_{n-1}z^{n}\|=\sup_{n\geq 1}\|F_{w}^{n}(1)\|=\infty.$
By the norm expression (4.2), we get the result in (ii).
(iii) The supercyclicity criterion with $D$ and $S$ as above, taking $n_{k}=k$, $k\geq 1$, is satisfied. Indeed, by Proposition 3.2, for each $\nu\geq 0$ the sequence $F_{w}^{*n}(k_{\nu})$ vanishes for large $n$.
(iv) By the assumptions in (4.3), $F_{w}$ becomes a bounded operator. Also, one can find $r<1$ and $N\in\mathbb{N}$ such that $\big{|}b_{n}/a_{n+1}\big{|}<r$ for every $n\geq N$. Hence, the conditions that appear in (i) and (ii) are easily seen to be equivalent.
Our next result on the chaos of $F_{w}^{*}$. In order to apply the chaoticity criterion, we need a representation of $f_{n}^{*}$ in terms of $k_{n}$. The proof is omitted as it runs parallel to the bilateral case, see Proposition 3.11.
@@
Let $k_{n}$ denote the co-ordinate functional $f\mapsto\frac{f^{(n)}(0)}{n!},$ defined on $\ell^{p}_{a,b}(\mathbb{D})$ or $c_{0,a,b}(\mathbb{D})$, where $n\geq 0$. We have
$f_{n}^{*}=\frac{1}{a_{n}}k_{n}+\sum_{j=1}^{n}(-1)^{j}\frac{b_{n-1}b_{n-2}\cdots b_{n-j}}{a_{n}a_{n-1}\cdots a_{n-j}}k_{n-j},$
@@
We now have necessary and sufficient conditions for $F_{w}^{*}$ to be chaotic on the dual of $\ell^{p}_{a,b}(\mathbb{D})$, stated below. We recall that, in [32] the authors provided a sufficient condition for the adjoint $F^{*}$ of the unweighted shift $F$ to be chaotic on a general analytic reproducing kernel space over $\mathbb{D}$ which includes the case of $\ell^{2}_{a,b}(\mathbb{D})$.
@@
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}<\infty.$
@@
$\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty.$
@@
To see that (ii) implies (iii), let $\psi=\sum_{m=0}^{\infty}\lambda_{m}f_{m}^{*}$ be a non-zero periodic vector for $F_{w}^{*}$ on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$, where $f_{m}^{*}$, $m\geq 0$, forms an unconditional Schauder basis for $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$. Now, using Proposition 4.5, we get
$\psi=\sum_{m=o}^{\infty}A_{m}k_{m},~{}~{}~{}~{}\text{where}~{}~{}~{}A_{m}=\lambda_{m}\frac{1}{a_{m}}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{m+j}\frac{b_{m}b_{m+1}\cdots b_{m+j-1}}{a_{m}a_{m+1}\cdots a_{m+j}},$
@@
for all $m\geq 0.$ Now, let $\nu\in\mathbb{N}$ be such that $(F_{w}^{*})^{\nu}\psi=\psi$. Then $(F_{w}^{*})^{n\nu}\psi=\psi$ for all $n\geq 1$. Equating the coefficient of $k_{0}$ both sides, we get
$\frac{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}{a_{n\nu}}\left(\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right)=\frac{1}{a_{0}}\left(\lambda_{0}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{j}\frac{b_{0}\cdots b_{j-1}}{a_{1}\cdots a_{j}}\right).$
Since the series in the right hand side converges, we get that
@@
$C\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}=\left\lvert a_{0}\right\rvert^{q}\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q},$
for some $C>0$. Note that using Hölder inequality, we can get
@@
$\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q}<\infty,$
@@
as $\{\lambda_{n}\}_{n}\in\ell^{q}(\mathbb{N})$ and $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty.$ Therefore,
@@
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}<\infty.$ Similarly, equating the coefficient of $k_{j},j=1,2,\cdots,\nu-1$, and performing calculations as above, it is possible to observe that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu+j}}{w_{n\nu+j-1}w_{n\nu+j-2}\cdots w_{j}}\right\rvert^{q}<\infty.$ We conclude that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}<\infty.$